Your house has SIX lights that are each used an average of four hours every day, and a 4.8 kW oven that is used an average of one hour a day.

Calculate the monthly energy consumption and waste heat produced for the following energy efficient appliances:

LED Bulbs 10W, 90% efficient

Energy Start Oven – 20% les energy, 95% efficient
Answer #1

Ans) Given,

6 light bulbs 10W each 90% efficient

4.8 kW Oven used 1 hr a day 95% efficient

Energy consumed by bulbs in a day = 6 x 10

= 60 W /day

Energy wasted = 0.1 x 60

= 6 W/day

Net energy consumed = 60 - 6 = 54 W

Energy consumed by oven in a day = 4800 W /day

Energy wasted per day by oven = 4800 x 0.05

= 240 W

Net energy consumed = 4800 - 240

= 4560 W

Total energy consumed in a day = 54 + 4560

= 4614 W

Total energy consumed in a month = 4614 x 30

= 138420 W or 13842 kW /month

or 138420 Joules

Total energy wasted in a day = 6 + 240 = 246 W

Total Energy wasted in a month = 246 x 30

= 7380 W/month

= 7380 Joules